# Empowering Scientific Workflows with Federated Agents

### v2

@@ 3.3.2. Agent
An Agent is a multithreaded entity that executes a behavior and manages communication with other entities. It is instantiated with a behavior, unique identifier (the address of the agent’s mailbox in the exchange), and exchange interface. An agent is a callable object that when run: (1) invokes the on_setup() callback of the behavior, (2) starts each @loop method in a separate thread, (3) spawns a thread to listen for new messages in the agent’s mailbox, and (4) waits for the agent to be shut down. An @action method is executed in a thread pool when requested remotely so as to not block the handling of other messages. Behaviors can optionally specify the maximum action concurrency.
Agents are designed to be long-running, but can be terminated by sending a shutdown request. Upon shutdown, the shutdown Event, passed to each @loop, is set; running threads are instructed to shutdown and waited on; and the on_shutdown() callback is invoked. Alternatively, an agent can terminate itself by setting the shutdown event. Similarly, an exception raised in an @loop method will shutdown the agent by default but can optionally be suppressed to keep the agent alive. Exceptions raised when executing @action methods are caught and returned to the remote caller.
The use of multi-threading means that behavior implementations must be aware of the caveats of Python’s global interpreter lock (GIL). Compute-heavy actions can dispatch work to other parallel executors, such as process pools, Dask Distributed (Rocklin, 2015), Parsl (Babuji et al., 2019), or Ray (Moritz et al., 2018). We discuss these patterns in more detail in Section 3.4.4. In the future, we would like to support async behaviors and exchanges for improved I/O performance, but scientific computing libraries in Python are not typically async compatible. In Python 3.13 and later, we provide experimental support for free-threading builds, which disable the GIL, enabling full multi-core performance. At this time, however, third-party library support for free-threading builds is limited.
Our decision to decouple behavior definitions from agent execution is deliberate. As behaviors encode application-level logic, we want them to be easily testable and reusable, independent of deployment details. Existing code bases can trivially transition an existing class definition into an agent by inheriting from Behavior and decorating with @action as needed, and behavior classes can still be used independently (i.e., not as a running agent).
3.3.3. Handles
Interacting with an agent is asynchronous; an entity sends a message to the agent’s mailbox and waits to receive a response message in its own mailbox. A handle is a client interface to a remote agent used to invoke actions, ping, and shutdown the agent. Each handle acts as a reference to that agent, translating each method call into a request message that is sent via the exchange and returning a Future. The handle also listens for response messages and accordingly sets the result on the appropriate Future. Rather than creating a return mailbox and listener thread for each handle that a client or agent may have, Academy will multiplex communication for multiple handles within a single process through a single mailbox. This multiplexing ensures that only one mailbox listener thread is needed per process (i.e., agent or client).
3.3.4. Exchange
Entities communicate by sending and receiving messages to and from mailboxes. An exchange hosts these mailboxes, and the Exchange protocol defines the interface to an exchange. Namely, the Exchange defines methods for registering new agent or client mailboxes, sending and receiving messages, and creating handles to remote agents. Registering an agent or client involves creating a unique ID for the entity, which is also the address of its mailbox, and initializing that mailbox within the exchange.
A mailbox has two states: open and closed. Open indicates that the entity is accepting messages, even if, for example, an agent has not yet started or is temporarily offline. Closed indicates permanent termination of the entity and will cause MailboxClosedError to be raised by subsequent send or receive operations to that mailbox.
Exchanges also provide mechanisms for agent discovery by querying based on agent behaviors. This also works with superclasses of behaviors. Consider, for example, a behavior ProteinFolder that can fold proteins (Anfinsen, 1973) and another behavior OpenProteinFolder that inherits from ProteinFolder and specifically uses the OpenFold model (Ahdritz et al., 2024). Querying for ProteinFolder would return the IDs of all agents inheriting from ProteinFolder whereas querying for OpenProteinFolder would return only specific agents using the OpenFold model. In the future, agents could provide additional metadata to enhance discovery.
Users can define custom exchanges to address specific hardware or application characteristics. We provide two exchange implementations for local and distributed agent deployments. The thread exchange stores messages in-memory and is suitable for agents running in separate threads of a single process, such as when testing.
The distributed exchange enables communication between entities across wide-area networks. Core to the distributed exchange is an object store that persists information about registered entities. A hybrid approach is used for message passing: direct messaging is preferred, and indirect message passing via the object store is available as a fallback. Upon startup, an entity writes its location (i.e., address and port) to the object store; peers that want to send a message can attempt to send directly to the entity’s address. If the peer is offline or a direct connection fails, such as in the presence of NAT (network address translation) or firewall restrictions, messages are appended to the list of pending messages in the object store. Entities continuously listen to incoming messages from peers and pending messages in the object store. Entities cache successful communication routes locally to reduce queries to the object store. Our implementation use TCP (transmission control protocol) sockets for direct messaging and a Redis server as the object store. Redis provides low-latency communication and optional replication, but applications that need greater fault-tolerance could consider DHT-based (distributed hash table) object stores.
We optimize the exchange for low latency, as control messages are typically small: $O$(100) bytes. However, action request and response messages can contain arbitrarily sized serialized values for arguments and results that can induce considerable overheads when messages are sent indirectly via the object store. To alleviate these overheads, we pass large values by reference and perform out-of-band data transfers by using ProxyStore (Pauloski et al., 2023, 2025), which provides pass-by-reference semantics in distributed computing through proxy objects. Proxy objects act like references (cheap to serialize and communicate) and automatically de-reference themselves to the true object using performant data storage and communication methods. For example, ProxyStore can leverage RDMA (remote direct memory access) transfers via Mochi (Ross et al., 2020) and UCX (Shamis et al., 2015), GridFTP via Globus Transfer (Chard et al., 2014), and reliable peer-to-peer UDP (user datagram protocol) through NAT hole-punching. ProxyStore also provides two key optimizations useful within Academy: proxies can be forwarded to actions executed on other agents without incurring additional data transfers and proxies can be asynchronously resolved to overlap communication and computation.
3.3.5. Launcher
An agent can be run manually, but the intended method of execution is via a launcher, which manages the initialization and execution of agents on remote resources. The Launcher protocol defines a launch() method with parameters for the behavior, exchange, and agent ID and returns a handle to the launched agent. Users can create custom implementations; we provide the following four that cover most use cases:
@@ 3.3.5. Launcher
  • Thread: Runs agents in separate threads of the same process. Useful for local development and testing or for light-weight or I/O bound agents.
  • Process: Runs agents in separate processes on same machine.
@@
  • Parsl: Runs agents across the workers of a Parsl Executor (Babuji et al., 2019). Parsl supports execution across local, remote, and batch compute systems.
@@
  • Globus Compute: Runs agents across Globus Compute Endpoints (Chard et al., 2020). Globus Compute is a cloud-managed function-as-a-service (FaaS) platform which can execute Python functions across federated compute systems.
The last three launchers support mechanisms to automatically restart agents if they exit unexpectedly. It is common for different agents in an application to be executed with different launchers, but all agents must be registered to the same exchange to interact.
1from academy.exchange.thread import ThreadExchange
2from academy.launcher.thread import ThreadLauncher
@@
5with Manager(
6 exchange=ThreadExchange(), # Can be swapped with
7 launcher=ThreadLauncher(), # other implementations
8) as manager:
9 behavior = Example() # From Listing 1
10 handle = manager.launch(behavior)
11
12 future = handle.square(2)
13 assert future.result() == 4
14
15 handle.shutdown() # Or via the manager
16 manager.shutdown(handle.agent_id, blocking=True)
@@
A Manager combines an exchange and one or more launchers to provide a single interface for launching, using, and managing agents. Each manager has a single mailbox in the exchange and multiplexes that mailbox across handles to all of the agents that it manages. This reduces boilerplate code, improves communication efficiency, and ensures stateful resources and threads are appropriately cleaned up. An end-to-end example is provided in 2.
3.4. Common Patterns
We have introduced basic building blocks necessary to build multi-agent systems and deploy agents across federated infrastructure. Now we discuss several common patterns that highlight features of Academy and guide users in building new agentic workflows.
3.4.1. State Checkpoints
Research infrastructure can fail; thus, agents may want to perform periodic state checkpointing. The framework does not enforce a specific checkpointing mechanism, as the format, location, and frequency of checkpoints are highly application specific, but on_startup() callbacks can be used to restore state automatically. For convenience, we provide a State API that provides a dictionary-like interface and persists values to the local file system.
3.4.2. Migration
Research infrastructure is typically static, so Academy does not require that the launcher provide mechanisms for automatic agent migration. Some launchers, such as Parsl, will restart agents on different workers if node-level failures cause agents to crash. Applications can also manually migrate agents across different launchers using agent shutdown and checkpointing mechanisms. These features are sufficient for users to implement custom launchers that enable automatic migration, such as to load-balance across resource pools.
3.4.3. Agent Hierarchies
Agents may dynamically need to create and manage child agents, either to offload tasks or to access new behaviors. A parent agent can create new child agents by using the same launcher used to create the parent, or by creating a new launcher. The use of different launchers is common in scenarios where parent agents want to initialize a local multi-agent system. For example, a client may launch an initial set of agents across federated resources using Globus Compute, and then those initial agents spawn more agents on local resources through Parsl.
3.4.4. Resource Pools
High-performance workflows may need to distribute work across many computers. In an agentic model, resource pool allocation can take two forms: agent managed resource pools or agents as resource pools. In the former, an agent allocates a pool of resources using a parallel computing framework, such as Parsl or Ray, and the agent’s actions dispatch work to resources in the pool. In the second pattern, we deploy identical agents across a set of resources and then route action requests across this agent pool (akin to worker pools in HTTP frameworks).
3.4.5. Process-as-a-Service
FaaS systems, such as Globus Compute, provide optimized execution of short-lived, stateless, and ephemeral functions. Academy agents can extend FaaS systems with process-as-a-service capabilities (Copik et al., 2022), enabling applications to utilize longer-lived, stateful, and isolated processes on-demand.
4. Evaluation
@@
We studied the performance characteristics of Academy to answer key questions including: How well does the system scale? How fast can agents be deployed? What is the messaging latency? We also make comparisons to Dask and Ray, two popular frameworks with support for distributed actors in Python. Although Academy agents provide a superset of features provided by actors, these evaluations contextualize the performance of the framework. In these comparisons, we use the terms agent and actor interchangeably.
@@
We conducted experiments using the Aurora supercomputer at the Argonne Leadership Computing Facility (ALCF), unless otherwise stated. Aurora has 10 6241062410\,62410 624 nodes interconnected by an HPE Slingshot 11 network and a high performance DAOS storage system. Aurora nodes contain two Intel Xeon Max CPUs, each with 52 physical cores and 64 GB of high-bandwidth memory; 512 GB of DDR5 memory per socket; and six 128 GB Intel Data Center Max GPUs. In some cases we also use the Polaris supercomputer at ALCF and the compute-zen-3 nodes of Chameleon Cloud’s CHI@TACC cluster (Keahey et al., 2020). Polaris has $560$ nodes interconnected by an HPE Slingshot 11 network. Polaris nodes contains one AMD EPYC Milan processor with 32 physical cores, 512 GB of DDR4 memory, and four 40 GB NVIDIA A100 GPUs. Each compute-zen-3 node contains two 64-core CPUs and 256 GB memory. Experiments were performed using Python 3.10, AutoGen 0.5.1, Dask 2025.2.0, Globus Compute 3.5.0, Parsl 2025.03.03, and Ray 2.43.0.
4.1. Weak Scaling
@@
We measure weak scaling performance from two aspects: agent startup and action completion time. The object store of the exchange is located on the head node of the Aurora batch job to best match the behavior of Dask and Ray.
@@
Figure 3. (Top) Warm-start time for $n$ agents/actors between Academy (using the Parsl launcher), Dask Actors, and Ray Actors. Ray does not benefit from warm-starts because a new process is spawned for each actor. (Bottom) Time to execute 30 actions per agent/actor (weak scaling). Each action sleeps for 1 s. Note the Academy and Ray lines are overlapped.
@@
[(Top) Warm-start time for $n$ actors/agents between Academy (using the Parsl launcher), Dask Actors, and Ray Actors. (Bottom) Weak scaling evaluation of action completion time.]
4.1.1. Agent Startup Time
We measure the time to spawn $n$ agents in Figure 3 (top). We pre-warm the worker processes by starting and stopping $n$ agents, then record the average startup time over five runs. Specifically, we measure the time between submitting the first agent to receiving a ping from all agents to ensure that they have finished their startup sequence. We configured Academy to use Parsl’s High-throughput Executor as the launcher. Ray always spawns a new process per actor and thus does not benefit from pre-warmed workers leading to high startup overheads at smaller scales. The cold start time with Academy and Dask is comparable to that of Ray and dominated by loading libraries from the shared file system. With warm starts, Academy starts a single actor in 5.5 ms, 2.8×2.8$\times$2.8 × faster than Dask. Academy scales well, starting $3328$ actors in 7.6 s compared to Dask’s 23.4 s, but Ray demonstrates an advantage at this scale with a 3.2 s startup. Since Academy can leverage many launcher types, applications requiring frequent startup of agents can utilize Parsl for low-latency, and applications launching thousands of long-running agents could use Ray.
4.1.2. Action Completion Time
In Figure 3 (bottom), we execute 30 sleep tasks (1 s) per agent and record the total completion time. We set the maximum concurrency to 1 for all agents to ensure that tasks are processed sequentially. Completion time remains constant for Academy and Ray up to 3328 agents while the performance of Dask degrades starting at 104 actors.
4.2. Academy Distributed Exchange
Next, we study the performance of the distributed exchange.
4.2.1. Data Transfer
Figure 4. (Top) Time for a client to invoke a no-op action on an actor as a function of input and output payload size with different optimizations enabled on the distributed exchange. Two scenarios are considered: client and agent are at the same site (left) and different sites (right). (Bottom) Time for a client to invoke a chain of $n$ actions across $n$ agents with a payload size of 10 MB. Each action in the chain is a no-op that passes the input data along to the next agent, and returns the resulting data. The pass-by-reference optimization reduces communication costs among intermediate actions.

### v3

@@ III-C1 Agent Runtime
Academy provides an agent Runtime that executes an agent and manages communication with other entities. It is instantiated with an agent class or instance, a unique identifier (the address of the agent’s mailbox in the exchange), and an exchange interface. When started, the Runtime: (1) invokes the agent_on_startup() callback of the agent, (2) spawns tasks for each @loop method, (3) spawns a task to listen for new messages in the agent’s mailbox, and (4) waits for the agent to be shut down. We use asyncio to ensure that many @action requests can be handled concurrently without blocking control loop and message listening tasks.
III-C2 Execution
An agent can be run manually, but the intended method of deployment is via an executor. Any class that implements the concurrent.futures.Executor protocol can be used to launch agents on distributed resources via the Manager interface (R1). We use the following executors which cover most local and remote resource types:
@@ III-C2 Execution
  • Thread: Runs agents in separate threads of the same process. Useful for local development and testing or for lightweight or I/O bound agents.
  • Process: Runs agents in distinct processes on a machine.
@@
  • Parsl: Runs agents across the workers of a Parsl Executor [32]. Parsl supports execution on local, remote, and batch compute systems.
@@
  • Globus Compute: Runs agents across Globus Compute Endpoints [15]. Globus Compute is a cloud-managed function-as-a-service (FaaS) platform which can execute Python functions across federated compute systems.
III-C3 State API
The Academy State API provides a dictionary-like interface for agents to persist in-memory state to storage. This API currently supports only writing/reading state to/from disk, but could be extended to other storage modalities. Agents can define the agent_on_startup() callback to restore state automatically. Given that research infrastructure can fail, agents may want to perform periodic state checkpointing (R5). Academy does not enforce a specific checkpointing mechanism, as the format, location, and frequency of checkpoints are highly application specific, agents can use the State API to periodically store checkpoints.
III-D Agent Communication
Entities communicate by sending and receiving messages to and from mailboxes. The mailboxes serve as mediated communication channels to decouple agents in time (R3); messages persist in a mailbox even when an agent is offline. An exchange hosts these mailboxes, and the Exchange protocol defines the interface to an exchange. Namely, the ExchangeClient defines methods for registering new agent or user mailboxes, sending and receiving messages, and creating handles to remote agents. Registering an agent or user involves creating a unique ID for the entity, which is also the address of its mailbox, and initializing that mailbox within the exchange.
A mailbox has two states: active and terminated. Active indicates that the entity is accepting messages, even if, for example, an agent has not yet started or is temporarily offline. Terminated indicates that the entity is permanently finished and will cause a terminated error to be raised by subsequent send or receive operations to that mailbox.
Users can define custom exchanges to address specific hardware or application characteristics (R2). We provide three exchange implementations: for testing (Thread Exchange), single-site (Hybrid), and distributed (Cloud) deployments.
III-D1 Thread Exchange
This implementation stores messages in-memory and is suitable for agents running in separate threads of a single process, such as when testing.
III-D2 Hybrid Exchange
This implementation enables communication between entities across local networks (e.g., HPC interconnects). It leverages an object store that persists information about registered entities. A hybrid approach is used for message passing: direct messaging is preferred, and indirect message passing via the object store is available as a fallback. Upon startup, an entity writes its location (i.e., address and port) to the object store; peers that want to send a message can attempt to send directly to the entity’s address. If the peer is offline or a direct connection fails, messages are appended to the list of pending messages in the object store. The multi-path design also makes the hybrid exchange suitable for dealing with heterogeneous networks.
Entities continuously listen to incoming messages from peers and pending messages in the object store. Entities cache successful communication routes locally to reduce queries to the object store. Our implementation uses TCP (Transmission Control Protocol) sockets for direct messaging and a Redis server as the object store. Redis provides low-latency communication and optional replication, but applications that need greater fault-tolerance could consider (distributed hash table (DHT)-based object stores. The security model for the hybrid exchange relies on Redis’s built-in password authentication and access control lists.
III-D3 Cloud Exchange
This implementation supports federated agent deployments. It exposes a secure, HTTP-accessible REST API that wraps Redis queues and enforces authorized access to mailboxes (R4).
The cloud exchange security model is designed to ensure that any communication on the exchange is done by authenticated and permitted agents. This guarantees, for instance, that any action request served by an agent comes from an authorized agent/user and similarly response data is only read by the requesting agent/user. Authorization and permissions enforcement occur at the exchange server; communication with the exchange is secured using standard TLS. Users are authenticated via Globus Auth, an identity and access management platform that supports federated authentication via thousands of supported identity providers [33]. Each agent is registered as an authenticated entity with the Globus Auth service. Creating an agent involves creating a new application identity with Globus Auth; a mailbox on the exchange linked with that identity, and a delegated token for the new agent to authenticate with the exchange; and launching the agent with the delegated token.
We assume that the launching entity is trusted, and that the launching mechanism is secure. When launched, the new agent exchanges a delegated token for a refresh token, allowing long-lived access to the exchange without reauthentication. The refresh token is used to obtain access tokens to the exchange; these are introspected by the exchange to verify the client ID matches the expected client of the mailbox, and are cached for 60 seconds. Revocation of the delegated token invalidates the refresh token and any derived access token, but the exchange does not respond to the revocation until the cache expires. This response time depends on the cache length. Users can revoke tokens using the Globus Python SDK or web app. Currently, permissions are coarse-grained—an agent is either allowed to communicate or not—but we plan to support finer-grained access controls for agents.
III-D4 Pass-by-Reference
We optimize the Hybrid and Cloud exchanges for low latency, as control messages are typically small: OO(100) bytes. However, action request and response messages can contain arbitrarily sized serialized values for arguments and results that can induce considerable overheads when messages are sent indirectly via the object store. To alleviate these overheads, we pass large values by reference and perform out-of-band data transfers by using ProxyStore [34, 35], which provides pass-by-reference semantics in distributed computing through proxy objects. Proxy objects act like references (cheap to serialize and communicate) and automatically de-reference themselves to the true object using performant data storage and communication methods. For example, ProxyStore can leverage RDMA (remote direct memory access) transfers via Mochi [36] and UCX [37], Globus Transfer [38], and reliable peer-to-peer UDP (user datagram protocol) through NAT hole-punching. Two key ProxyStore optimizations are useful within Academy: proxies can be forwarded to actions executed on other agents without incurring additional data transfers and proxies can be asynchronously resolved to overlap communication and computation.
1from concurrent.futures import ThreadPoolExecutor
2from academy.exchange.local import LocalExchangeFactory
@@ III-D4 Pass-by-Reference
5async with await Manager.from_exchange_factory(
6 # Can be swapped with other implementations
7 factory=LocalExchangeFactory(),
8 executors=ThreadPoolExecutor(),
9) as manager:
10 agent = Example() # From Listing 1
11 handle = await manager.launch(agent)
12
13 result = await handle.square(2)
14 assert result == 4
15
16 await handle.shutdown() # Or via the manager
17 await manager.shutdown(handle, blocking=True)
@@
IV Evaluation
@@
We studied the performance characteristics of Academy to answer specific questions: How well does the system scale? How fast can agents be deployed? What is the messaging latency? In non-federated settings, we also compare to Dask and Ray, two popular frameworks with support for distributed actors in Python.
@@
We conducted experiments using the Aurora supercomputer at the Argonne Leadership Computing Facility (ALCF), unless otherwise stated. Aurora has 10 62410\,624 nodes interconnected by an HPE Slingshot 11 network and a high performance DAOS storage system. Aurora nodes contain two Intel Xeon Max CPUs, each with 52 physical cores and 64 GB of high-bandwidth memory; 512 GB of DDR5 memory per socket; and six 128 GB Intel Data Center Max GPUs (split into two tiles, or logical GPUs each). In some cases we also use the Polaris supercomputer at ALCF and the compute-zen-3 nodes of Chameleon Cloud’s CHI@TACC cluster [39]. Polaris has $560$ nodes interconnected by an HPE Slingshot 11 network. Polaris nodes contains one AMD EPYC Milan processor with 32 physical cores, 512 GB of DDR4 memory, and four 40 GB NVIDIA A100 GPUs. Each compute-zen-3 node contains two 64-core CPUs and 256 GB memory. Experiments were performed using Python 3.10, AutoGen 0.5.1, Dask 2025.2.0, Globus Compute 3.5.0, Parsl 2025.03.03, and Ray 2.43.0.
IV-A Weak Scaling
@@
We measure weak scaling performance from two aspects: agent startup and action completion time. We use the hybrid exchange with the object store located on the head node of the batch job to best match the behavior of Dask and Ray.
@@
Figure 3: (Top) Warm-start times for $n$ agents/actors with Academy (with Parsl launcher), Dask Actors, and Ray Actors. Ray does not benefit from warm starts because a new process is spawned for each actor. (Bottom) Time to execute 30 actions per agent/actor (weak scaling). Each action sleeps for 1 s. Note the Academy and Ray lines are overlapped.
IV-A1 Agent Startup Time
@@
IV-A2 Action Completion Time
In Figure 3 (bottom), we execute 30 sleep tasks (1 s) per agent and record the total completion time. We set the maximum concurrency to 1 for all agents to ensure that tasks are processed sequentially. Completion time remains constant for Academy and Ray up to 3328 agents while Dask performance degrades starting at 104 actors.
IV-B Academy Exchange
Figure 4: Time for a client to invoke a no-op action on an actor as a function of input and output payload size with different optimizations enabled on the hybrid and cloud exchanges. Two scenarios are considered: client and agent are at the same site using the hybrid exchange (left) and different sites using the cloud exchange (right).
Next, we investigate performance and optimizations of the hybrid and cloud exchange implementations. In baseline, all message data are communicated indirectly between peers via the exchange’s object store. For node to node communication on Aurora (Aurora → Aurora), the object store of the hybrid exchange is located on the head node of the Aurora batch job. For remote communication, (Workstation → Aurora), the cloud exchange is deployed on AWS. In pass-by-ref, messages are still communicated via the object store, but action arguments and results are replaced with references using ProxyStore. ProxyStore is configured to use ZeroMQ and P2P endpoints for intra-site and inter-site transfer of referenced objects, respectively. In direct, messages are communicated directly between peers, circumventing the object store; this is only possible when peers are located within the same site.
In Figure 4, we measure the time it takes for a client to invoke a no-op action on an agent as a function of payload size. We compare baseline, pass-by-ref, and direct across two scenarios: Aurora → Aurora, where the client and the agent are located on different Aurora nodes and messages are passed via the hybrid exchange, and Workstation → Aurora, where the client is located on a personal workstation, the agent is on an Aurora node, and messages are passed via the cloud exchange. The round trip network latencies are: Aurora compute node to Aurora head node: 0.23 ms; Aurora to AWS: 8.8 ms; Aurora to Workstation: 2.2 ms; and Workstation to AWS: 10.4 ms.
We observe that network latency to the exchange limits performance at smaller payload sizes ($\leq$ 100 KB). Direct, which is possible only in the intra-site scenario, circumvents these latencies. In both scenarios, pass-by-ref alleviates overheads of data transfer to and from the exchange by communicating data directly between the client and agent via ProxyStore. For intra-site transfers, pass-by-ref and direct reduce action latency compared to the baseline by 91.2% and 97.6%, respectively, with 100 MB payloads. For inter-site transfers, pass-by-ref reduces action latency by 45.1%.
IV-C Agent Messaging
Here, we investigate the performance of agent messaging.
Figure 5: (Top) No-op action latency between two agents/actors running on separate Aurora nodes versus action input and output payload size. (Middle) Action latency between agents located on a workstation and Aurora. Academy is compared to Globus Compute, a tool for building federated workflows that only support stateless functions, so state is read from the file-system on every invocation. (Bottom) Completion time for a simulated two agent chat where agents send ten messages back and forth with varied message sizes. Academy is compared to AutoGen’s distributed runtime.
IV-C1 Action Latency